OpenStudy (anonymous):

use logarithm properties to expand each question:

OpenStudy (anonymous):

log(a^2b^3/c^5)

OpenStudy (whpalmer4):

There are two properties of logarithms which will be quite useful here: \[\log (ab) = \log(a) + \log(b)\]\[\log (c/d) = \log(c) -\log(d)\] With those, you can rewrite your equation as the sum of 3 logarithms. Each of those may be further simplified by using \[\log x^n = n \log x\]